# Federico Echenique and Farzad Pourbabaee

### 14 Jan 2024 v1

@@ 3.1 Optimality and Equilibrium
Let $\mathcal{E}$ be a convex exchange economy, with no aggregate uncertainty, and an aggregate endowment of $\omega=\bm{1}$. Assume $f$ is a weakly Pareto optimal allocation. For any given $\varepsilon>0$, let $\mathcal{V}^{(\varepsilon)}\coloneqq\sum_{i\in I}\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ be the Minkowski sum of the $\varepsilon$-approximate upper contour sets. Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Then
@@
Figure 2 illustrates the concepts in the statement of Theorem 2. The figure depicts a case with two states, $s_{1}$ and $s_{2}$, and two individual traders. For pictorial simplicity we take $\varepsilon=0$ in the graphs. Aggregate consumption is $\omega$, and entails no aggregate uncertainty. The square defined by the origin and $\omega$ is an Edgeworth box that describes all the allocations between the two agents. The Pareto optimal allocations are on the solid curve that connects the origin and $\omega$, while $f=(f_{1},f_{2})$ is a particular Pareto optimal allocation (indicated by the tangency of the two agents’ indifference curves at the point $f$, with the usual Edgeworth-box convention that one agent’s coordinate system has been rotated $180^{\circ}$ and its origin coincides with $\omega$). For the allocation $f$ of the initial endowment $\omega$, we may consider all the aggregate consumptions that can be decentralized among the two agents so as to obtain higher utility than in $f$. This set of aggregate bundles, the Scitovsky contour $\mathcal{V}^{(0)}$, consists of all the bundles to the north-east of the dashed curve passing through $\omega$ (a curve that, at $\omega$, has the same slope as the common slope of each individual agent’s utility at $f$). [^5] The Scitovsky contour is a key concept in the proof of the second welfare theorem. For further discussion of this concept, see Debreu (1951) and Samuelson (1956).
@@
Debreu considers the minimum amount of aggregate resources (call it $\omega^{*}$) that could be used to provide agents with the same utility as in $f$, and thinks of the gap between $\omega$ and $\omega^{*}$ as the inefficiency inherent in the allocation $f$. In Debreu’s words, these are “nonutilized resources.” He proposes to measure this gap by means of a “distance with economic meaning:” $p\cdot(\omega-\omega^{*})$, where $p$ is an “intrinsic price vector” associated with $\omega^{*}$. To obtain a measure that is, in a sense, scale independent, he works with the ratio of $p\cdot\omega^{*}$ to $p\cdot\omega$. Prices $p$ follow from an argument that is analogous to the second welfare theorem; in particular, they are not uniquely defined.
@@
Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Under the hypotheses of Theorem 2, if $f$ is not weakly Pareto optimal, and $\mathrm{CRU}(f)$ its coefficient of resource utilization, then
@@ 3.2 Prior Beliefs and Welfare-Improving Trade
Yaari (1969) defines the subjective belief as a probability distribution vector that supports the upper contour set. When the upper contour has a kink (due, for example, to the ambiguity in preferences and its induced lack of differentiability), there will be multiple supporting vectors. In the spirit of Rigotti et al. (2008), we define the subjective belief set as the set of all supporting vectors. Specifically, let $f$ be an act in $\mathbb{R}_{+}^{d}$. The upper contour set of agent $i$ is $\{g:g\succeq_{i}f\}$, and the subjective belief set at $f$ is defined by
@@
The set $\mathcal{B}_{i}(f)$ is a convex and closed (hence compact) subset of the $d$-dimensional probability simplex $\Delta S$. One may also interpret a vector in $\mathcal{B}_{i}(f)$ as the set of normalized prices that support the consumption of the act $f$ [^6] The price interpretation is suitable for the exercise in Section 3.1, where we could have phrased our results in terms of the Walrasian theory of general equilibrium where there is no uncertainty and $S$ captures the space of available consumption goods. Here we are more tightly following the story of uncertain consumption in financial markets.
Following our notation from before, we define the $\delta$-extension of the subjective belief set by
@@
Let $\mathcal{E}$ be an exchange economy with preferences $\succeq_{i}\in\mathcal{P}$ for all $i\in I$ and no aggregate uncertainty. If the allocation $f$ is $\varepsilon$-Pareto dominated, then there exists a constant $c>0$, such that for every proper subset $J\subset I$,
@@
In particular, if in a two-agent economy, an allocation is not approximately Pareto optimal, then the volume of the subjective belief set of at least one agent is exponentially smaller than the volume of the probability simplex. Put differently, if there is a possibility for a strong welfare-improving trade, then the subjective belief set of at least one agent must be “very small.”
We interpret the theorem as saying that some degree of ambiguity neutrality is needed for the existence of a welfare-improving trade. That agents have a small set of prior beliefs seems to capture the idea of being ambiguity neutral, but the volume measure does not have an obvious behavioral counterpart. We seek to provide such a behavioral interpretation in Section 4.
@@ 4.1 Behavioral Implications of Theorem 3
In this section we study the behavioral implications of Theorem 3. For the purposes of this discussion, we consider a setting with two agents. The message of our theorem is that, if the current risk-sharing arrangement between the two agents is strongly Pareto dominated (namely if there exists a possibility of a strongly welfare improving ex-ante trade), then the volume of the underlying belief set of at least one agent must be very small. The smaller is this volume, the closer is that agent to the ambiguity neutrality — a connection we seek to quantify in this section.
It is well-established that ambiguity aversion leads to less trade. Theorem 3 offers a quantitative expression of this result: that the possibility of an $\varepsilon$-Pareto improving trade necessitates small ambiguity aversion in high dimensions.
This interpretation is in line with the comparative notion of ambiguity aversion proposed by Ghirardato et al. (2004). Specifically, they show in the max-min setting of Gilboa and Schmeidler (1989), the ambiguity aversion of a decision maker decreases as her multiple prior set shrinks with respect to the set inclusion order.
To establish the connection between ambiguity aversion and the volume of the belief set, we appeal to the max-min setting of Gilboa and Schmeidler (1989), in which there is a convex compact set of priors $\Pi\subseteq\Delta S$ and the agent’s cardinal evaluation of an act $f\in\mathbb{R}_{+}^{d}$ is $u(f)=\min_{\mu\in\Pi}f\,\cdot\,\mu$. We further assume that $\Pi$ has constant width, namely the distance between any distinct parallel supporting hyperplanes of $\Pi$ (residing on the $d$ dimensional probability simplex) is constant.
@@
When $\Pi$ has constant width, $\theta(f)$ becomes constant in $f$, and we write $\theta(f)\equiv\theta$. Thus we can take $\theta$ as a measure of ambiguity aversion. In the next proposition we show how the upper bound on the relative volume in Theorem 3 means that $\theta$ vanishes as $d\to\infty$.
@@ 5.1 Brunn-Minkowski Inequality
5.2 Proof of Results in Section 3.1
@@
Because of the union bound and (3.1) one has
@@
Since the volume is translation invariant, we can shift the subsets in the above inequality by $f_{i}$ and thus achieve the bound in (3.4) and thereby (3.3).∎
@@ 5.3 Concentration and Isoperimetric Inequalities
5.4 Proof of Results in Section 3.2

### 14 Feb 2024 v2

@@ 3.2 Optimality and resource utilization
Let $\mathcal{E}\colon I\to\mathcal{C}\times\mathbb{R}^{d}_{+}$ be a convex exchange economy, with no aggregate uncertainty, and aggregate endowment $\omega=\bm{1}$.
Assume $f$ is a weakly Pareto optimal allocation. For any given $\varepsilon>0$, let $\mathcal{V}^{(\varepsilon)}\coloneqq\sum_{i\in I}\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ be the Minkowski sum of the $\varepsilon$-approximate upper contour sets. Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Then
@@
Figure 2 illustrates the concepts in the statement of Theorem 2. The figure depicts a case with two states, $s_{1}$ and $s_{2}$, and two individual traders. For pictorial simplicity we take $\varepsilon=0$ in the graphs. Aggregate consumption is $\omega=(1,1)$, and entails no aggregate uncertainty. The square defined by the origin and $\omega$ is an Edgeworth box that describes all the allocations between the two agents. The Pareto optimal allocations are on the solid curve that connects the origin and $\omega$, while $f=(f_{1},f_{2})$ is a particular Pareto optimal allocation (indicated by the tangency of the two agents’ indifference curves at the point $f$, with the usual Edgeworth-box convention that one agent’s coordinate system has been rotated $180^{\circ}$ and its origin coincides with $\omega$). For the allocation $f$ of the initial endowment $\omega$, we may consider all the aggregate consumptions that can be decentralized among the two agents so as to obtain higher utility than in $f$. This set of aggregate bundles, the Scitovsky contour $\mathcal{V}^{(0)}$, consists of all the bundles to the north-east of the dashed curve passing through $\omega$ (a curve that, at $\omega$, has the same slope as the common slope of each individual agent’s utility at $f$). [^6] The Scitovsky contour is a key concept in the proof of the second welfare theorem, where it plays a role that is analogous to our use in the proof of Theorem 2. For further discussion of Scitovsky contours and its connection to the Kaldor criterion and cost-benefit analysis see Debreu (1951), Samuelson (1956), or Echenique (2021).
@@
It is also worth emphasizing that the basic message of Theorem 2 holds, even if we restrict attention to perturbations that result in a more expensive aggregate bundle. For Figure 2, the probability of the dotted region once we condition on a more expensive perturbation is quite close to one. This probability, however, converges exponentially to zero in $d$. A similar point holds for Figure 1 in the Introduction.
@@
Debreu considers a minimal amount of aggregate resources (call it $\omega^{*}$) that could be used to provide agents with the same utility as in $f$, and thinks of the gap between $\omega$ and $\omega^{*}$ as the inefficiency inherent in the allocation $f$. In Debreu’s words, these are “non-utilized resources.” He proposes to measure this gap by means of a “distance with economic meaning:” $p\cdot(\omega-\omega^{*})$, where $p$ is an “intrinsic price vector” associated with $\omega^{*}$. To obtain a measure that is, in a sense, scale independent, he works with the ratio of $p\cdot\omega^{*}$ to $p\cdot\omega$. Prices $p$ follow from an argument that is analogous to the second welfare theorem; in particular, they are not uniquely defined.
@@
Let $\mathcal{E}\colon I\to\mathcal{C}\times\mathbb{R}^{d}_{+}$ be an exchange economy under the the hypotheses of Theorem 2.
Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. If $f$ is not weakly Pareto optimal, and $\mathrm{CRU}(f)$ its coefficient of resource utilization, then
@@ 3.3 Prior Beliefs and Welfare-Improving Trade
Yaari (1969) defines a subjective belief as a probability distribution that supports an agent’s upper contour set at an act. When the upper contour has a kink (due, for example, to the ambiguity in preferences and its induced lack of differentiability), there will be multiple supporting vectors, and thus multiple beliefs (or multiple priors). In the spirit of Rigotti et al. (2008), we define the subjective belief set as the set of all supporting vectors.
Specifically, let $f$ be an act in $\mathbb{R}_{+}^{d}$. The upper contour set of agent $i$ is $\{g:g\succeq_{i}f\}$, and the subjective belief set at $f$ is defined by
@@
The set $\mathcal{B}_{i}(f)$ is a convex and closed (hence compact) subset of the $d$-dimensional probability simplex $\Delta S$. One may also interpret a vector in $\mathcal{B}_{i}(f)$ as the set of normalized prices that support the consumption of the act $f$ [^7] The price interpretation is suitable for the exercise in Section 3.1, where we could have phrased our results in terms of the Walrasian theory of general equilibrium when there is no uncertainty and $S$ captures the set of available consumption goods. Here we are more tightly following the story of uncertain consumption in financial markets.
We define the $\delta$-extension of the subjective belief set by
@@
Our first result adapts the characterization of Pareto optimality in Rigotti et al. (2008) to our setting, allowing for approximate Pareto domination.
@@
Let $\mathcal{E}\colon I\to\mathcal{P}\times\mathsf{E}^{d}_{+}$ be an exchange economy with no aggregate uncertainty.
If the allocation $f$ is $\varepsilon$-Pareto dominated, then there exists a constant $c>0$, such that for every proper subset $J\subset I$,
@@
The message of the theorem is clearest in a two-agent economy. The existence of a $\varepsilon$-Pareto improving trade is then only possible when one of the two agents has a small set of subjective beliefs, as measured through its Euclidean volume. We expand on this interpretation in Section 4.1.
More generally, in an economy with many agents, Theorem 3 says that, no matter how we partition the set of agents in two, at least one of the groups of agents must have diverse set of priors. The volume of the intersection of their sets of priors must be small. If we think of the intersection of the set of priors as corresponding to a representative agent for the group (a notion that may be formalized), then this agent must be close to being ambiguity neutral: see again the discussion in Section 4.1 for a sufficient condition under which the small volume conclusion leads to a behavioral property that bounds the degree of ambiguity aversion.
@@ 4.1 Behavioral Implications of Theorem 3
In this section we study the behavioral implications of Theorem 3. For the purposes of this discussion, we consider a setting with two agents. The message of our theorem is that, if the current risk-sharing arrangement between the two agents is strongly Pareto dominated (namely if there exists a possibility of a strongly welfare improving ex-ante trade), then the volume of the underlying belief set of at least one agent must be very small. The smaller is this volume, the closer is that agent to being ambiguity neutral — a connection we seek to quantify in this section. [^8] For an overview of ambiguity and ambiguity aversion, see Machina and Siniscalchi (2014) and Gilboa and Marinacci (2016).
It is well known that ambiguity aversion leads to less trade (see, for example, Dow and da Costa Werlang (1992)). Theorem 3 offers a quantitative expression of this fact: that the possibility of an $\varepsilon$-Pareto improving trade necessitates small ambiguity aversion in high dimensions.
Our interpretation of the size of a set of priors is in line with the comparative notion of ambiguity aversion proposed by Ghirardato et al. (2004). Specifically, they show in the max-min expected utility setting of Gilboa and Schmeidler (1989), that the ambiguity aversion of a decision maker decreases as her multiple prior set shrinks with respect to the set inclusion order.
To establish the connection between ambiguity aversion and the volume of the decision maker’s set of beliefs, or priors, we again appeal to the max-min expected utility setting of Gilboa and Schmeidler (1989). In their model, there is a convex and compact set of priors $\Pi\subseteq\Delta S$, and the agent’s choices over acts are guided by a utility function of the form
$u(f)=\min_{\mu\in\Pi}f\,\cdot\,\mu,$
for each act $f\in\mathbb{R}_{+}^{d}$.
We further assume that $\Pi$ has constant width, namely that the distance between any distinct parallel supporting hyperplanes of $\Pi$ (residing on the $d$ dimensional probability simplex) is constant. This assumption is crucial for our use of the result of Schramm (1988), but one can imagine a quantitative relaxation of the assumption of constant width and that it would translate into a modified bound on the degree of ambiguity aversion.
@@
When $\Pi$ has constant width, $\theta(f)$ becomes constant in $f$, and we write $\theta(f)\equiv\theta$. Thus we can take $\theta$ as a measure of ambiguity aversion. Its meaning is the difference in how an ambiguity averse and ambiguity loving agent would evaluate any given normalized act. A small degree of ambiguity aversion corresponds to a small difference between the ambiguity averse and ambiguity loving evaluation of an act.
In the next proposition we show how the upper bound on the relative volume in Theorem 3 means that $\theta$ vanishes as $d\to\infty$.
@@ 5.1 Brunn-Minkowski Inequality
5.2 Proofs of the results in Section 3.1
@@
The union bound and (3.1) yields
@@
Given that Euclidean volume is translation invariant, we can shift the measured sets in the above inequality by $f_{i}$, and thus achieve the bound in (3.4), and thereby (3.3).∎
@@ 5.3 Concentration and Isoperimetric Inequalities
5.4 Proofs of the results in Section 3.3